OpenStudy (anonymous):

Evaluate (5 [ 8 1/3 + 27 1/3 ]3 ) ^¼ all the steps please

OpenStudy (anonymous):

\[[5(8^{1/3} + 27^{1/3})^3]^{1/4}\]

OpenStudy (anonymous):

\(=[5(2+3)^3]^{1/4}=[5(5^3)]^{1/4}=(5^4)^{1/4}=5\).

OpenStudy (anonymous):

u deserve thousands of medals